19 & 20
019 (part 1 of 2) 10.0 points On the way to the moon, the Apollo astro- nauts reach a point where the Moons gravi- tational pull is stronger than that of Earths. Find the distance of this point from the center of the Earth. The masses of the Earth and the Moon are 5.98 x 1024 kg and 7.36x 1022 kg, respectively, and the distance from the Earth to the Moon is 3.84 x 108 m Answer im units of m. 020 (part 2 of 2) 10.0 points What would the acceleration of the astro- naut be due to the Earths gravity at this point if the moon was not there? The value of the universal gravitational constant is 6.672 x 10 11 N m2/kg2 Answer in units of m/s
Answer #1

Given

Me = 5.98 x 1024 kg

R = 3.84 x 108m

Mm = 7.36 x 1022 kg

Assume if re is the distance from this point to the center of the earth and rm is the distance from this point to the center of the moon

G*m*Me/(re2) = G*m*Mm/(rm2)……eq1

and

we know

re + rm = R……………eq2

solving for re = distance from center on the earth

G*m*Me/(re2) = G*m*Mm/(rm2)

Me/(re2) = Mm/(rm2)

rm / re = \sqrt(Mm/Me)

rm / re = \sqrt(7.36 x 1022/5.98 x 1024)

rm / re = 0.1109400392

replacing value for eq 2

re = R/(1+0.1109400392)

re = (3 x 108)/(1+0.1109400392)

re = 2.7 x 108 m

(20)

g' = G*Me/re2

g' = (6.672x10-11*5.98x1024)/(2.7 x 108)2

g' = 0.0054713 m/s2

g' = 5.47 x 10-3 m/s2